# Sequential Randomization Tests Using e-values: Applications for trial monitoring

### v4

@@ Appendix A R Code
median_stop <- median(first_crossing, na.rm = TRUE)
[…]
if (sim_type == "Power(Alternative)") {
[…]
return(list(rate = rate, median_stop = median_stop, n_patients = n_patients))
# Plot Trajectories
plot_trajectories_surv <- function(n_trials = 20,
n_patients = 400,
HR_true = 0.7,
alpha = 0.05,
title = "e-SurvivalTrajectories") {
trajectories <- list()
for (i in 1:n_trials) {
@@
labs(
title = title,
subtitle = sprintf("TrueHR=%.2f", HR_true),
x = "NumberofEvents",
y = "e-value(logscale)"
@@
# Run Scenarios
if (interactive()) {
# Define design parameters
target_hr <- 0.80
# 1. Check Type I Error (Null: HR = 1.0)
# Passing n_patients = NULL triggers the auto-calculation based on target_HR
simulate_eSurvival(
n_sims = 2000,
n_patients = NULL, # Auto-calculate N for 80% power
HR_true = 1.0, # Reality: Null
target_HR = target_hr,
alpha = 0.05
# 2. Check Power (Alternative: HR = 0.7)
simulate_eSurvival(
n_sims = 2000,
n_patients = NULL, # Auto-calculate N for 80% power
HR_true = 0.8, # Reality: Alternative works
target_HR = target_hr,
alpha = 0.05
# 3. Plot Trajectories
# We calculate N explicitly here just to pass it to the plot function
req_events <- ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(target_hr))^2)
p <- plot_trajectories_surv(
n_trials = 30,
n_patients = req_events,
HR_true = 0.8,
alpha = 0.05,
title = sprintf("e-Survival(TargetHR=%.2f,N=%d)", target_hr, req_events)
print(p)
set.seed(2)
p1<-plot_trajectories_surv(
n_patients = ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(0.8))^2),
HR_true = 0.8)
ggsave("traj_survivaleRT_08.pdf",p1,
width = 8, height = 5)
p2<-plot_trajectories_surv(
n_patients = ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(0.8))^2),
HR_true = 1)
p2
ggsave("traj_survivaleRT_null.pdf",p2,
width = 8, height = 5)
# APPENDIX: Verification of Staggered vs. Simultaneous Entry
# This script demonstrates that the "Simultaneous Entry" assumption used in the
# main simulation yields statistically identical results to a realistic
# "Staggered Entry" design, provided the analysis is based on "Time on Study".
verify_staggered_equivalence <- function(n_sims = 1000,
n_patients = 631, # Updated to match HR=0.8 scenario
HR_true = 0.8, # Updated to match HR=0.8 scenario
recruit_period = 12, # Recruitment over 12 months
@@
# 1. Generate the underlying biology (Survival Times)
# -- Method A: Simultaneous (The Simple Way) --
@@
# -- Method B: Staggered (The "Real World" Way) --
# We generate fresh biology for the staggered arm to compare distributions
@@
# The event happens at this Calendar Time:
calendar_event_time <- entry_time + trial_stag\$time
# The analyst calculates "Time on Study" (Analysis Time):
# Analysis Time = Calendar Event Time - Entry Time
calculated_study_time <- calendar_event_time - entry_time
# Run e-process on the calculated study time
# Note: compute_eSurvival handles the sorting of events internally
res_stag <- compute_eSurvival(calculated_study_time,
trial_stag\$status,
@@
Method = rep(c("Simultaneous(Simplified)", "Staggered(Realistic)"), each = n_sims)
# Summary Stats
print(df_res %>%
group_by(Method) %>%
summarise(
Mean_E = mean(e_value),
Median_E = median(e_value),
Power_Estimated = mean(e_value >= 1/alpha)
# 3. Visualization
p <- ggplot(df_res, aes(x = e_value, fill = Method)) +
@@
labs(
title = "EquivalenceofSimultaneousvs.StaggeredEntry",
subtitle = sprintf("Distributionoffinale-values(N=%d,HR=%.2f)", n_patients, HR_true),
x = "Finale-value",
y = "Density"
@@
return(p)
# Run the verification
if (interactive()) {
# Matching the N=631 / HR=0.8 scenario from the paper
p_verify <- verify_staggered_equivalence(
n_sims = 1000,
n_patients = 631,
HR_true = 0.8
print(p_verify)
# Optional: Save for supplement
ggsave("supp_staggered_equivalence.pdf", p_verify, width = 8, height = 5)

### v5

@@ Appendix A R Code
list(rate = rate, se = se, median_stop = median(first_crossing, na.rm = TRUE),
n_patients = n_patients)
plot_trajectories_surv <- function(n_trials = 20, n_patients = 400, HR_true = 0.7,
alpha = 0.05, title = "e-SurvivalTrajectories") {
trajectories <- lapply(1:n_trials, function(i) {
@@
labs(title = title, subtitle = sprintf("TrueHR=%.2f", HR_true),
x = "NumberofEvents", y = "e-value(logscale)") +
@@
# --- Staggered vs Simultaneous Entry Verification ---
verify_staggered_equivalence <- function(n_sims = 1000, n_patients = 631,
HR_true = 0.8, recruit_period = 12,
@@
# Simultaneous
@@
# Staggered
@@
calculated_study_time <- entry_time + trial_stag\$time - entry_time
res_stag <- compute_eSurvival(calculated_study_time, trial_stag\$status,
@@
Method = rep(c("Simultaneous", "Staggered"), each = n_sims)
ggplot(df_res, aes(x = e_value, fill = Method)) +
@@
labs(title = "EquivalenceofSimultaneousvs.StaggeredEntry",
subtitle = sprintf("N=%d,HR=%.2f", n_patients, HR_true),
x = "Finale-value", y = "Density") +
@@
# --- e-RTms: Multi-State Models ---
# States: 1=Ward, 2=ICU, 3=Home (absorbing), 4=Dead (absorbing)
P_ctrl <- matrix(c(
0.880, 0.070, 0.030, 0.020,
0.070, 0.915, 0.000, 0.015,
0.000, 0.000, 1.000, 0.000,
0.000, 0.000, 0.000, 1.000
), nrow = 4, byrow = TRUE)
P_trt <- matrix(c(
0.870, 0.050, 0.050, 0.030,
0.090, 0.900, 0.000, 0.010,
0.000, 0.000, 1.000, 0.000,
0.000, 0.000, 0.000, 1.000
), nrow = 4, byrow = TRUE)
BURN_IN <- 30
RAMP <- 50
MAX_DAYS <- 28
START_STATE <- 2
THRESHOLD <- 20
GOOD_TRANS <- c("2_1", "1_3")
simulate_patient_ms <- function(P, max_days = MAX_DAYS, start_state = START_STATE) {
state <- start_state
transitions <- NULL
for (day in 1:max_days) {
if (state %in% c(3, 4)) break
new_state <- sample(1:4, 1, prob = P[state, ])
if (new_state != state) {
transitions <- rbind(transitions, c(from = state, to = new_state, day = day))
state <- new_state
list(final_state = state, transitions = transitions)
simulate_trial_ms <- function(n_patients, P_trt, P_ctrl) {
all_trans <- NULL
all_arms <- c()
final_states <- data.frame(patient = 1:n_patients, arm = NA, final_state = NA)
for (i in 1:n_patients) {
arm <- sample(0:1, 1)
P <- if (arm == 1) P_trt else P_ctrl
pat <- simulate_patient_ms(P)
final_states\$arm[i] <- arm
final_states\$final_state[i] <- pat\$final_state
if (!is.null(pat\$transitions)) {
for (j in 1:nrow(pat\$transitions)) {
all_trans <- rbind(all_trans, pat\$transitions[j, ])
all_arms <- c(all_arms, arm)
list(transitions = all_trans, arms = all_arms, final_states = final_states)
compute_eRTms <- function(all_transitions, all_arms,
burn_in = BURN_IN, ramp = RAMP,
good_trans = GOOD_TRANS) {
n <- nrow(all_transitions)
if (is.null(n) $|$ n == 0) return(1)
wealth <- numeric(n)
wealth[1] <- 1
n_good_trt <- 0; n_total_trt <- 0
n_good_ctrl <- 0; n_total_ctrl <- 0
for (i in 1:n) {
trans <- all_transitions[i, ]
arm <- all_arms[i]
trans_key <- paste0(trans["from"], "_", trans["to"])
is_good <- trans_key %in% good_trans
if (i > burn_in && n_total_trt > 0 && n_total_ctrl > 0) {
c_i <- min(1, max(0, (i - burn_in) / ramp))
rate_trt <- n_good_trt / n_total_trt
rate_ctrl <- n_good_ctrl / n_total_ctrl
delta <- rate_trt - rate_ctrl
lambda <- if (is_good) 0.5 + 0.5 * c_i * delta else 0.5 - 0.5 * c_i * delta
} else {
lambda <- 0.5
lambda <- max(0.01, min(0.99, lambda))
mult <- if (arm == 1) lambda / 0.5 else (1 - lambda) / 0.5
wealth[i] <- if (i > 1) wealth[i - 1] * mult else mult
if (arm == 1) {
n_total_trt <- n_total_trt + 1
if (is_good) n_good_trt <- n_good_trt + 1
} else {
n_total_ctrl <- n_total_ctrl + 1
if (is_good) n_good_ctrl <- n_good_ctrl + 1
return(wealth)
run_simulation_ms <- function(n_sims, n_patients, P_trt, P_ctrl,
threshold = THRESHOLD, verbose = TRUE) {
rejections <- 0
first_crossing <- numeric(n_sims)
n_transitions <- numeric(n_sims)
if (verbose) pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
for (s in 1:n_sims) {
trial <- simulate_trial_ms(n_patients, P_trt, P_ctrl)
if (is.null(trial\$transitions) $|$ nrow(trial\$transitions) < BURN_IN) {
first_crossing[s] <- NA
n_transitions[s] <- 0
next
wealth <- compute_eRTms(trial\$transitions, trial\$arms)
n_transitions[s] <- length(wealth)
crossing <- which(wealth >= threshold)
if (length(crossing) > 0) {
rejections <- rejections + 1
first_crossing[s] <- crossing[1]
} else {
first_crossing[s] <- NA
if (verbose) setTxtProgressBar(pb, s)
if (verbose) close(pb)
rate <- rejections / n_sims
se <- sqrt(rate * (1 - rate) / n_sims)
list(
rate = rate,
se = se,
median_crossing = median(first_crossing, na.rm = TRUE),
median_transitions = median(n_transitions)
plot_trajectories_ms <- function(n_trials = 30, n_patients, P_trt, P_ctrl,
threshold = THRESHOLD, title = "") {
trajectories <- list()
for (i in 1:n_trials) {
trial <- simulate_trial_ms(n_patients, P_trt, P_ctrl)
if (!is.null(trial\$transitions) && nrow(trial\$transitions) > BURN_IN) {
wealth <- compute_eRTms(trial\$transitions, trial\$arms)
trajectories[[length(trajectories) + 1]] <- data.frame(
transition = 1:length(wealth), wealth = wealth, trial = i
df <- bind_rows(trajectories)
ggplot(df, aes(x = transition, y = wealth, group = trial)) +
geom_line(alpha = 0.4, color = "steelblue") +
geom_hline(yintercept = threshold, linetype = "dashed", color = "red") +
geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
scale_y_log10() +
labs(title = title, x = "TransitionNumber", y = "e-value(logscale)") +
annotate("text", x = Inf, y = threshold * 1.5,
label = sprintf("1/alpha=%.0f", threshold),
color = "red", hjust = 1.1) +
theme_minimal() +
theme(plot.title = element_text(face = "bold"))